The Lakeside Coffee Shop morning customer load follows a normal distribution. The mean number of customers is 60 and the standard deviation is 15. If the number of customers falls below 42, then two staff members are sufficient. What are the chances only two workers are required tomorrow? Include evidence of your work by typing out your solution fully.


1
Expert's answer
2021-08-05T13:36:27-0400

Let "X=" the number of customers: "X\sim N(\mu, \sigma^2)."

Given "\mu=60, \sigma=15."


"P(X<42)=P(Z< \dfrac{42-\mu}{\sigma})"

"=P(Z< \dfrac{42-60}{15})=P(Z<-1.2)"

"\approx0.115070"

The probability that only two workers are required tomorrow is 0.115070.